# Exotic Superfluid with Emergent Flux in a One-Dimensional Bose-Fermi Mixture

### v3

@@ Abstract
Intriguingly, the central charge obtained from entanglement entropy in DMRG calculations remains consistently $\approx 2$, as illustrated in Fig. 4. This observation indicates that, despite the projected particle’s contribution to optimizing the system’s total energy, it resembles the original particles. Global excitations are contributed by bosons superfluid and fermions Luttinger liquid. It is worth mentioning that the condensation of a bosonic superfluid at a momentum $k\neq 0$ has been previously explored in various works, including those studying systems with correlated hopping Struck et al. (2012); Keilmann et al. (2010), as well as in the experimental realization of a two-dimensional topological chiral superfluid observed in optical lattice experiments Wang et al. (2021); Kock et al. (2015); Soltan-Panahi et al. (2012). Nevertheless, the microscopic origin of the one-dimensional CSF in our system is fundamentally different. It arises directly from the interplay between two species of fermions and serves as a background field to optimize the system’s overall energy.
CSF in Extended BF Hubbard Model.— We have elucidated the origin of the 1D CSF phase using a model that incorporates the kinetic energy of composite fermions. However, this seemingly “artificial” term is not an indispensable component. The 1D CSF phase may naturally arise in the extended BF Hubbard model, which is given by
$H_{BF}=-\sum_{\langle ij\rangle}(t_{f}c^{\dagger}_{i}c_{j}+t_{b}b^{\dagger}_{i}b_{j}+H.c.)+U_{bf}\sum_{i}n^{b}_{i}n^{f}_{i}$
$+\frac{U_{bb}}{2}\sum_{i}n^{b}_{i}(n^{b}_{i}-1)+\frac{V_{bf}}{2}\sum_{\langle ij\rangle}(n^{b}_{i}n^{f}_{j}+n^{f}_{i}n^{b}_{j}).$ (2)
where $t_{b}$ represents the hopping amplitude of bosons, while $t_{f},U_{bb},n^{b}_{i},n^{f}_{i}$ retain the same meanings as in Hamiltonian (1).
Here we introduce an asymmetry by setting an unequal number of bosons and fermions. This asymmetry leads to strong scattering of composite particles by surplus fermions, resulting in instability. The Hamiltonian incorporates a standard on-site boson-fermion repulsion term $U_{bf}$ollet et al. (2006), facilitating the formation of composite particles. To prevent phase separation, a nearest-neighbor attraction $V_{bf}$ is introduced. The filling factor of fermion is less than 1/3 to avoid collapse stateRizzi and Imambekov (2008) and $N_{f}\approx 2N_{b}$. When $t_{f}\approx 10t_{b}$ and $U_{bf}$ is one order larger than $t_{f}$ ($U_{bf}=U_{bb}$), DMRG simulations show that bosons condensate at $|k_{0}|=2\pi/L$ with even fermion number $N_{f}$, as shown in Fig. 5(c) similar to that in Fig. 5(a). For odd $N_{f}$ in Fig. 5(d), bosons condensate at $k=0$ similar to that in Fig. 5(b). The fermions parity-dependent chirality provides further confirmation, which suggests that this novel chirality can be controlled by simply adding or removing particles through tuning chemical potentials. The realization of the CSF phase is more feasible in practical experimental settings.
@@
Experimental feasibility.— With the thriving development of quantum simulation experimentsBloch et al. (2008); Gross and Bloch (2017); Schfer et al. (2020); Weitenberg and Simonet (2021), our model is promising to be effectively simulated in an optical lattice utilizing a combination of ultracold bosonic and fermionic atomsTruscott et al. (2001); Stan et al. (2004); Inouye et al. (2004); Fukuhara et al. (2009). Experimentally, periodic boundary conditions have been achieved in a 1D ring-shaped optical lattice with finite sitesFranke-Arnold et al. (2007); Amico et al. (2005); Houston et al. (2008); Amico et al. (2014, 2022). The characteristic signature of the CSF phase can be observed through widely utilized time-of-flight (TOF) experiments, which restore phase coherence information throughout the entire latticeGreiner et al. (2002); Bloch et al. (2008). The density distribution captured in TOF images corresponds to the momentum distribution of the trapped atoms. During TOF expansion, the fermionic atoms influenced by an effective flux induced by the bosonic CSF will exhibit a doughnut-shaped pattern (analogous to the right panel of Fig.5 in Ref.Amico et al. (2022)). This pattern contrasts with the bell-shaped pattern observed when bosons are in a normal superfluid state and fermions experience no effective flux Amico et al. (2022).
@@
Conclusions.— Our study reveals a novel many-body effect in a 1D BFM system, where the CSF phase emerges as the ground state and spontaneously breaks TSR to optimize the energy of the whole system. Concrete numerical evidence and theoretical analysis provide a comprehensive understanding of the origin of CSF. Notably, the ground state can be tuned between the CSF and a conventional superfluid by adjusting chemical potentials. Furthermore, we demonstrate the experimental feasibility of realizing this innovative state. Our discovery opens up opportunities to explore new quantum phases of matter.
@@
We thank Ting-Kuo Lee and Kai Chang for their fruitful suggestions. This work is supported by the National Key Research and Development Program of China Grant No. 2022YFA1404204, and the National Natural Science Foundation of China Grant No. 12274086.

### v4

@@ Abstract
Intriguingly, the central charge obtained from entanglement entropy in DMRG calculations remains consistently $\approx 2$, as illustrated in Fig. 5. This observation indicates that, despite the projected particle’s contribution to optimizing the system’s total energy, it resembles the original particles. Global excitations are contributed by bosons superfluid and fermions Luttinger liquid. It is worth mentioning that the condensation of a bosonic superfluid at a nonzero momentum has been previously explored in theoretical works, including those studying systems with correlated hopping Struck et al. (2012); Keilmann et al. (2010), as well as in the experimental realization of a two-dimensional topological chiral superfluid in optical lattices Wang et al. (2021); Kock et al. (2015); Soltan-Panahi et al. (2012). Nevertheless, the microscopic origin of the one-dimensional CSF in our system is fundamentally different. It arises directly from the interplay between two species of fermions and serves as a background field to optimize the system’s overall energy.
@@
Discussions.— With the rapid advancement of quantum simulation techniques Bloch et al. (2008); Gross and Bloch (2017); Schfer et al. (2020), a diverse array of interacting BFMs has been experimentally explored Stan et al. (2004); Milczewski and Duda (2023); Duda et al. (2023); Truscott et al. (2001); Fukuhara et al. (2009); Inouye et al. (2004). Our system, characterized by a density imbalance between the two species, can be readily realized by loading unequal numbers of ultracold bosonic and fermionic atoms into a 1D optical lattice. Tunable Feshbach resonances provide precise control over the short-range boson-fermion interaction strength Chin et al. (2010); Stan et al. (2004); Inouye et al. (2004). Employing light fermions and heavy bosons, such as 6Li-23NaStan et al. (2004) and 6Li-41KLiu et al. (2018); Wang et al. (2020), naturally results in an inequality between the tunneling amplitudes of the two species. Furthermore, the tunnelling amplitude can be finely tuned by adjusting the depth of the species-dependent lattice potentialLiu et al. (2004); Soltan-Panahi et al. (2011). Experimentally, periodic boundary conditions have been achieved in a 1D ring-shaped optical lattice with finite sitesFranke-Arnold et al. (2007); Amico et al. (2005); Houston et al. (2008); Amico et al. (2014, 2022); Chetcuti et al. (2023). The characteristic signature of the CSF phase can be observed through widely utilized time-of-flight (TOF) experiments, which restore phase coherence information throughout the entire latticeGreiner et al. (2002); Bloch et al. (2008). The density distribution captured in TOF images corresponds to the momentum distribution of the trapped atoms. During TOF expansion, the fermionic atoms influenced by an effective flux induced by the bosonic CSF will exhibit a doughnut-shaped pattern (analogous to the right panel of Fig.5 in Ref.Amico et al. (2022) and Fig.4 in Ref.Chetcuti et al. (2023)). This pattern contrasts with the bell-shaped pattern observed when bosons are in a normal superfluid state and fermions experience no effective flux Amico et al. (2022).
@@
Our study reveals a novel many-body effect in a 1D BFM system, where the CSF phase emerges as the ground state and spontaneously breaks TSR to optimize the energy of the whole system. Concrete numerical evidence and theoretical analysis provide a comprehensive understanding of the origin of CSF. Notably, the ground state can be tuned between the CSF and a conventional superfluid by adjusting chemical potentials. Furthermore, we demonstrate the experimental feasibility of realizing this innovative state. Our discovery opens up opportunities to explore new quantum phases of matter.
Acknowledgements.
@@
Acknowledgements.— We thank Ting-Kuo Lee and Kai Chang for their fruitful suggestions. This work is supported by the National Key Research and Development Program of China Grant No. 2022YFA1404204, and the National Natural Science Foundation of China Grant No. 12274086.